Solve the equation dy/dx= -(x2+y2)/xy


1
Expert's answer
2021-08-05T16:28:17-0400

Let us solve the equation "\frac{dy}{dx}= -\frac{x^2+y^2}{xy}" using the transformation "y=ux." Then "\frac{dy}{dx}=\frac{du}{dx}x+u," and hence we get the equation "\frac{du}{dx}x+u= -\frac{1+u^2}{u}," which is equivalent to "\frac{du}{dx}x= -\frac{1+u^2}{u}-u," and hence to "\frac{du}{dx}x= -\frac{1+2u^2}{u}."


It follows that


"\frac{udu}{1+2u^2}= -\frac{dx}{x}"


"\int\frac{udu}{1+2u^2}= -\int\frac{dx}{x}"


"\frac{1}{4}\int\frac{4udu}{1+2u^2}= -\int\frac{dx}{x}"


"\int\frac{d(1+2u^2)}{1+2u^2}= -4\int\frac{dx}{x}"


"\ln(1+2u^2)= -4\ln|x|+\ln|C|"


"\ln(1+2u^2)+4\ln|x|=\ln|C|"


"\ln(x^4(1+2u^2))=\ln|C|"


"x^4(1+2u^2)=C"


"x^4(1+2(\frac{y}{x})^2)=C"


We conclude that the general solution of the equation "\frac{dy}{dx}= -\frac{x^2+y^2}{xy}" is


"x^4+2x^2y^2=C"